OpenStudy (anonymous):

(m-n)^2-(2m+3n)^2 Factor - the answer is -(m+4n)(3m+2n) Don't know how to get there?

OpenStudy (campbell_st):

well look at it this way (m - n)^2 = m^2 - 2mn + n^2 - (2m+3n)^2 = 4m^2 + 12mn + 9n^2 do the subtraction ------------------- -3m^2 -14mn - 8n^2 factor out the negative -(3m^2 + 14mn + 8m^2) now factor there quadratic... start with 3m^2 3 is prime so use 3m and m all terms are positive so then so it becomes -(3m + ?? n)(m + ? n) so you just need the factors of 8 so the middle term becomes 14mn